Assume that the average annual salary for a worker in the United States is $40,000 and that the annual salaries for Americans are normally distributed with a standard deviation equal to $6,500. Find the following:

(A) What percentage of Americans earn below $35,000?
(B) What percentage of Americans earn above $45,000?

Answer by stanbon(75887)

Assume that the average annual salary for a worker in the United States is $40,000 and that the annual salaries for Americans are that the annual salaries for Americans are normally distributed with a standard deviation equal to $6,500. Find the following:
(A) What percentage of Americans earn below $35,000?
z(35000) = (35000-40000)/6500 = -0.7692
P(x< 35000) = P(z< -0.7592) = 0.2209 = 22.09%
==============================================
(B) What percentage of Americans earn above $45,000?
z(45000) = (45000-40000)/6500 = +0.7692
P(x> 45000) = P(z > 0.7692 = 0.2209 = 22.09%
=================================================
Cheers,
Stan H.